OpenStudy (anonymous):

the ratio of males to females in a room is 5 to 4. if 5 males and 6 femaes leave the room, the ratio of males to females changes to 4 to 3. how many people are in the room initially?

OpenStudy (anonymous):

that's better.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh cmon it is not that bad i bet. smart minds that we have here

myininaya (myininaya):

m/f=5/4 (m-5)/(f-6)=4/3

OpenStudy (anonymous):

\[\frac{m}{f}=\frac{5}{4}\] \[4f=5f\] \[\frac{m-5}{f-6}=\frac{4}{3}\] \[3(m-5)=4(f-6)\]

OpenStudy (anonymous):

oops i meant \[4m=5f\] "typo"

OpenStudy (anonymous):

\[m=\frac{5}{4}f\]

OpenStudy (anonymous):

the answer is supposed to be 81 i dont know how that's supposed to work

OpenStudy (anonymous):

we're getting there... slowly

OpenStudy (anonymous):

ohhh ok

myininaya (myininaya):

\[4m=5f\] \[m=\frac{5f}{4}\] and we have \[\frac{m-5}{f-6}=\frac{4}{3}\] \[3(m-5)=4(f-6)\] \[3m-15=4f-24\] but m=5f/4 \[3*\frac{5f}{4}-15=4f-24\] \[\frac{15f}{4}-15=4f-24\] multiply both sides by 4 \[15f-60=16f-96\] \[-60+96=16f-15f\] \[36=1f=f\]

myininaya (myininaya):

so \[m=\frac{5*f}{4}=\frac{5*36}{4}=5*9=45\]

myininaya (myininaya):

f+m=36+45=81

myininaya (myininaya):

81 peeps are in the room 36 are women 45 are men

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

\[3(\frac{5}{4}f-5)=4(f-6)\] \[\frac{15}{4}f-15=4f-24\] \[9=\frac{1}{4}f\] \[f=36\]

OpenStudy (anonymous):

yes did you get 36 woment?

myininaya (myininaya):

satellite you are too slow lol

OpenStudy (anonymous):

old and slow it is true but i wanted to solve with pencil before i wrote something stupid

myininaya (myininaya):

by the way if you didn't realize i called f females and m males but i think that was obvious

OpenStudy (anonymous):

oh really because i had it reversed

myininaya (myininaya):

lol

myininaya (myininaya):

maybe the men in your room had too much estrogen in there system

OpenStudy (anonymous):

they were fmen and mwomen that is all. makes for trouble in the restrooms at the bar

myininaya (myininaya):

lol

myininaya (myininaya):

he was kidding we called f females/women and m males/men